OpenStudy (amenah8):

find dy/dx using logarithmic differentiation: y = x^cosx CLICK to see my work so far...

OpenStudy (amenah8):

lny=lnx^cosx --> lny=cosxlnx --> (1/y)(dy/dx)=(lnx)(-sinx)

OpenStudy (anonymous):

\[\ln y=\cos x~\ln x\] \[\frac{ 1 }{ y }\frac{ dy }{ dx }=\cos x*\frac{ 1 }{ x }+\left( -\sin x \right)\ln x\] \[\frac{ dy }{ dx }=?\]

OpenStudy (amenah8):

what is (1/y)(dy/dx)? (dy/dyx)?

OpenStudy (anonymous):

\[\frac{ dy }{ dx }=y \left( \frac{ \cos x }{ x }- \sin x~\ln x \right)\] substitute the value of y

OpenStudy (amenah8):

and that would be my final answer, because i cannot simplify anymore!

OpenStudy (amenah8):

thank you so much!

OpenStudy (superdavesuper):

@Amenah8 wait wait wait - did u substitute the formula for y back in?

OpenStudy (amenah8):

when he said to substitute the value for y, did he mean to put in y for something else?

OpenStudy (superdavesuper):

@surjithayer suggested that u should substitute y = x^cosx so that dy/dx is expressed only in x

OpenStudy (amenah8):

so put into x^cosx for y?

OpenStudy (amenah8):

(dy/dx) = (x^cosx))(cosx/x)-sinxlnx) ?

OpenStudy (superdavesuper):

correct :) sorry about my late reply!

OpenStudy (amenah8):

it's okay! that's my final answer, right?

OpenStudy (superdavesuper):

well i THINK so...but dont kill me if im wrong! lol

OpenStudy (amenah8):

haha thank you!

OpenStudy (superdavesuper):

welcome